OpenStudy (amina_soneviseth):

Medal and fan 7 = 3/5t -2

OpenStudy (michele_laino):

is your equation like this: \[7 = \frac{3}{5}t - 2\]

OpenStudy (amina_soneviseth):

yes, @Michele_Laino

OpenStudy (anonymous):

easy problem if so

OpenStudy (anonymous):

t=15

OpenStudy (anonymous):

add 2 to both sides

OpenStudy (amina_soneviseth):

no direct answers please

OpenStudy (michele_laino):

If I add +2 to both sides I can write this: \[7 + 2 = \frac{3}{5}t - 2 + 2\] please simplify

OpenStudy (amina_soneviseth):

so it would be 9 = 3/5

OpenStudy (texaschic101):

don't forget your t

OpenStudy (michele_laino):

ok! Now I myultiply both sides by 5: \[9 \cdot 5 = \frac{3}{5}t \cdot 5\] please simplify again

OpenStudy (amina_soneviseth):

kk! :D

OpenStudy (texaschic101):

or you can divide both sides by 3/5

OpenStudy (amina_soneviseth):

45 = 3t

OpenStudy (anonymous):

then divide both sides by 3

OpenStudy (michele_laino):

correct! Finally I divide both sides by 3: \[\frac{{45}}{3} = \frac{{3t}}{3}\]

OpenStudy (anonymous):

do i get a medal too? I helped

OpenStudy (amina_soneviseth):

yes! ill fan u @themazinmathmaster

OpenStudy (amina_soneviseth):

@Michele_Laino thank for the help!

OpenStudy (anonymous):

the full answer with steps: http://www.imagebam.com/image/5f1657441116169